OpenStudy (anonymous):

The curve defined by x^3+xy-y^2=10 has a vertical tangent line when x=? How does one go about solving this here problem?

OpenStudy (anonymous):

implicit differentiation, solve for dy/dx = infinity (or the denominator is zero.)

OpenStudy (anonymous):

3x^2+x(dy/dx)+y=2y(dy/dx)=10 ?

OpenStudy (anonymous):

* -2y(dy/dx)

OpenStudy (anonymous):

good, except it should equal zero

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so -( 3x^2 +y) /(x-2y)

OpenStudy (anonymous):

wait a minute, how did you get the divided?

OpenStudy (anonymous):

(Can also find dx/dy and set equal to 0.)

OpenStudy (anonymous):

nvm, I get it, what do you do after that?

OpenStudy (calculusfunctions):

\[x ^{3}+xy -y ^{2}=10\] Now implicitly differentiate to get\[3x ^{2}+y +x \frac{ dy }{ dx }+2y \frac{ dy }{ dx }=0\]Now solve this equation for dy/dx. Thus\[\frac{ dy }{ dx }=\frac{ -3x ^{2}-y }{ x +2y }\]For a vertical tangent, the derivative is undefined. This implies that\[x +2y =0\] \[x =-2y\] Substitute x into the given equation. \[(-2y)^{2}+(-2y)(y)-y ^{2}=10\] \[4y ^{2}-2y ^{2}-y ^{2}=10\] \[y ^{2}=10\] \[y =\pm \sqrt{10}\] Therefore the vertical tangents are at \[x =\pm \sqrt{10}\]

OpenStudy (anonymous):

I think you're a bit off @calculusfunctions

OpenStudy (anonymous):

The given eauation is x^3, not 3x^2

OpenStudy (anonymous):

that and x=2y

OpenStudy (anonymous):

not x^2, rather

OpenStudy (calculusfunctions):

I know that! Look at my work. I said that the derivative is 3x²

OpenStudy (calculusfunctions):

Oh yes that is true! That was a careless mistake. The correct answer is \[x =\pm2\sqrt{10}\]

OpenStudy (anonymous):

how would you get that answer if x is cubes,a nd cannot simplify with the y^2?

OpenStudy (anonymous):

would you not use the quadratic formula?

OpenStudy (anonymous):

@calculusfunctions, my good man, that is not the correct answer.

OpenStudy (anonymous):

web assign ? or solution key?

OpenStudy (anonymous):

Worksheet, A. 0 or -1/3 B. 1.037 C. 2.074 D. 2.096 E. 2.154

OpenStudy (anonymous):

your teacher is a d*** it's 2.074

OpenStudy (anonymous):

Many apologies, but I do not know how you would come to that answer. Would you care to explain?

OpenStudy (anonymous):

just what he did, but correctly.

OpenStudy (anonymous):

(haha) Yes, but I came up with -1/4 would you then use that to solve for x?

OpenStudy (anonymous):

rather, -1/8

OpenStudy (anonymous):

no. y is 1.0371 and x is 2y

OpenStudy (anonymous):

how did you find y...

OpenStudy (anonymous):

plug in y=2x (it's a cubic) find the root.

OpenStudy (anonymous):

wouldn't y=x/2?

OpenStudy (anonymous):

whoops x=2y plug in x=2y (it's a cubic) find the root.

OpenStudy (anonymous):

either way though should work.

OpenStudy (anonymous):

8y^3+y^2=10, right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

would you then use the quadratic formula?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

it's not a quadratic

OpenStudy (anonymous):

you use the cubatic formula aka Wolfram

OpenStudy (anonymous):

I was solving it this way and it occurred to me there must be some simpler way your teacher wants you to do the problem... but I haven't figured that bit out yet.

OpenStudy (calculusfunctions):

OK! I see my mistake. So give me a few minutes to type it up again.

OpenStudy (calculusfunctions):

\[x ^{3}+xy -y ^{2}=10\] \[3x ^{2}+y +x \frac{ dy }{ dx }-2y \frac{ dy }{ dx }=0\] \[\frac{ dy }{ dx }=\frac{ 3x ^{2}+y }{ 2y -x }\]Now \[2y -x =0\] \[x =2y\] Substitute x = 2y into the given equation. \[(2y)^{3}+(2y)(y)-y ^{2}=10\] \[4y ^{2}+2y ^{2}-y ^{2}=10\] \[5y ^{2}=10\] \[y ^{2}=2\] \[y =\pm \sqrt{2}\] \[ x =\pm2\sqrt{2}\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

it should be 2.07424

OpenStudy (anonymous):

you're 0-2 tonight @calculusfunctions

OpenStudy (anonymous):

Thanks algebraic, I looked at the directions and I could use my Graphing calc, so I just did that. Thanks again, to both of you. :)

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

that's why it was a goofy answer.

OpenStudy (anonymous):

:3 sorry

OpenStudy (anonymous):

np:)

OpenStudy (calculusfunctions):

@Algebraic! How is it 2.07424? Why don't you offer your solution?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Um, y should be 1.0371208412063, and x=2y, sooo... x=2.07424

OpenStudy (anonymous):

Turns out you can't do this just by hand... ?

OpenStudy (anonymous):

Again, my good man, it is x^3, not x^2

OpenStudy (calculusfunctions):

@Jack, that is x³, but when you type it using the equation feature at the bottom, it looks like x². If you don't believe me, why don't you try doing x³ yourself, using the equation feature at the bottom left.

OpenStudy (anonymous):

4y2+2y2−y2=10 5y2=10 y2=2 y=±2√ x=±22√ this is what you typed, copy paste, good sir.

OpenStudy (anonymous):

(2y)3+(2y)(y)−y2=10 4y2+2y2−y2=10 5y2=10 y2=2 y=±2√ x=±22√ This is all of it, so yes, I see there WAS an x^3, but you changed it to x^2

OpenStudy (calculusfunctions):

And I did get x = 2y. My good man, you need to look again to see that I have x = 2y the second time I typed the solution.

OpenStudy (calculusfunctions):

Of course it changes to 3x squared when you take the derivative of x cubed.

OpenStudy (calculusfunctions):

My second solution is therefore perfect. I'm sorry you don't see that.

OpenStudy (anonymous):

Many apologies, but your answer may still be incorrect, my good man.

OpenStudy (anonymous):

THIS is the second solution you allowed me: Substitute x = 2y into the given equation. (2y)3+(2y)(y)−y2=10 4y2+2y2−y2=10 5y2=10 y2=2 y=±2√ x=±22√

OpenStudy (anonymous):

A simply typo, which resulted in an incorrect calculation.

OpenStudy (calculusfunctions):

Ah I see when I plugged x = 2y back into the original equation, I did (2y)²... instead of (2y)³. In that case you are right. I apologize.

OpenStudy (anonymous):

No apologies are required, my good man. :)

OpenStudy (calculusfunctions):

Thanks, I appreciate that but I was wrong nonetheless.

OpenStudy (calculusfunctions):

Be more careful typing next time.

OpenStudy (anonymous):

Good man! This was a good challenge, nonetheless. Quite fun, was it not? Lessons learned: always check work, and that lesson applies to the best of us, does it not?

OpenStudy (anonymous):

Thank you for your time and effort, it is greatly appreciated.

OpenStudy (calculusfunctions):

Of course. I agree with everything you said. Thank you!